# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 7 Towards Non-adaptive Group Testing
Section 3 concerned adaptive group testing. In this section, we study the SNAGT model which is a model we proposed in Section 2 as a model in between adaptive and non-adaptive testing. To recall, in SNAGT, tests are designed beforehand and do not depend on prior test results. They are done sequentially and we know the result of a test just after performing it. While the test design does not depend on the result of the previous tests, we assume that we can stop testing at any point. In the following, we first focus on the SNAGT model, propose a testing strategy and analyze its performance. Then we discuss challenges one faces in designing fully non-adaptive algorithms.
As we already saw, when the testing is done adaptively, we get the best results when the number of infections is concentrated around its mean $\mu$. In other words, with high probability $|e^{*}|=\mu+o(\mu)$. In this section, we consider the same assumption, i.e., the size of the infected set is concentrated around its mean.
The following theorem shows that using $O(\mu H(X)+\mu\log n)$ tests, in the SNAGT mode, we can find the target edge with high probability.
@@
Suppose for a graph $G$, the probability distribution $\mathcal{D}$ is such that $\Pr_{e\sim\mathcal{D}}(|\left|e\right|-\mu|>\epsilon\mu)=o(1)$ for any constant $\epsilon>0$. There is an algorithm that performs $O(\mu\cdot H(X)+\mu\log n)$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
@@
To simplify the proof here, we assume that every edge has the same size $d$. A very similar and more general proof is given in Appendix E.
We follow a proof technique similar to [LCH+14, Theorem 2]. In particular, we partition the edges into subgraphs so that the edge probabilities in the same subgraph are close to each other. Suppose the edges are partitioned into subgraphs $G_{1},G_{2},\ldots,G_{k}$ where $G_{i}=(V,E_{i})$ and $\forall e\in E_{i}:\frac{1}{2^{i}}\leq p(e)\leq\frac{1}{2^{i-1}}$. We can assume $k\leq n\log n$ if we want to have an error that goes to zero, as any edge in $e_{i},i>n\log n$ has probability at most $\frac{1}{2^{n\log n-1}}$ and there are at most $2^{n}$ of them, so the total error probability from this is at most $\frac{2^{n}}{2^{n\log n-1}}\rightarrow 0$. Denote $m_{i}=|E_{i}|$ as the number of edges in $G_{i}$, and $\mathbb{P}(G_{i})=\sum_{e\in E_{i}}p(e)$ the probability that the target edge is sampled from $G_{i}$.
First, imagine that we have the information that the target edge is in the $i$’th hypergraph. We perform the tests randomly, i.e. we pick each node with probability $1/d$ independently. Then, using [GLS22, Theorem 4], we can find the target edge with high probability using $d\log m_{i}$ tests. So on average, the number of tests is:
$\mathbb{E}[T]$ $=\sum_{i}\mathbb{P}(G_{i})d\log(m_{i})$
$=d\sum_{i}(\sum_{e\in G_{i}}p(e))\log(m_{i}).$
But $m_{i}\leq 2^{i}$ and $1/2^{i}\leq p(e)\leq 1/2^{i-1}$, so $m_{i}\leq 2/p(e)$ for all $e\in E_{i}$. Hence we get
$\mathbb{E}[T]$ $=d\sum_{i}(\sum_{e\in G_{i}}p(e))\log(m_{i})$
$\leq d\sum_{i}\sum_{e\in G_{i}}p(e)\log(2/p(e))$
$=d\sum_{e\in E}p(e)(1+\log(1/p(e)))$
$=d+dH(X).$
So in this case, we get $O(dH(X))$ tests in expectation.
But of course, we do not know which $G_{i}$ contains the target edge. To fix this, we propose the following algorithm: keep testing randomly as before, and rule out inconsistent edges in each subgraph until at least one subgraph has exactly one edge. Let $SG$ be the set of subgraphs with exactly one edge at this point. Then perform $2d\log n$ additional tests one by one and rule out inconsistent edges. If the single edge remains in exactly one subgraph after the tests, return the edge as the target edge. If there is more than one subgraph in $SG$ with exactly one edge, return one of them at random. If at one point during the $2d\log n$ tests, single edges are ruled out in all subgraphs of $SG$, continue the random test until another component has exactly one edge again and repeat. We proved that conditioned on the target edge being in the $i$’th subgraph, the algorithm performs $O(d\log m_{i}+d\log n)$ tests. Hence by the above argument, we perform $O(dH(X)+d\log n)$ tests in total.
Now we prove that the error probability goes to zero. Suppose the target edge is in $i$’th subgraph. We know that if the algorithm performs $O(d\log m_{i})$ tests, there is only one edge remaining in it with high probability and after the $2d\log n$ “sanity check” tests, it remains in the graph. Let $err_{1}$ be the event that the algorithm outputs an edge before performing $d\log m_{i}$ tests. Then it means the algorithm found one or more subgraphs with a single edge in them, and after performing $d\log n$ tests, these subgraphs are not ruled out. The probability that a non-target edge survives after $2d\log n$ tests is $1/n^{2}$ and there are at most $n$ of them, as the number of subgraphs is bounded by $n$ and the edge is the only one remaining in the subgraph. Hence by union bound, this probability is at most $\mathbb{P}(err_{1})<1/n$. Also, define $err_{2}$ be the event that after $O(d\log m_{i})$ tests, more than one edge remains in $G_{i}$. In [GLS22], they showed that this has a probability that goes to zero. Hence, the total probability of error goes to zero.
We have not been able to generalize the above technique to get general upper bounds for non-adaptive group testing. To recall, in non-adaptive testing, all the tests are performed at once. In this case, each test is independent of the previous ones, and only after seeing the results of all tests can we recover the infected set. Non-adaptive group testing has been extensively studied in the literature for independent group testing [LCH+14, WGG23, PR08], and typically, state of the art non-adaptive methods are differing from the adaptive counterpart only by $O(\log n)$ or $O(d)$ (multiplicative factor). For instance, $O(d^{2}\log n)$[DHH00] is needed in the classic non-adaptive combinatorial setting (compared to $O(d\log n)$ tests for adaptive) or $O(\log n(H(X)+\mu))$[LCH+14] in non-adaptive independent group testing (compared to $O(H(X)+\mu)$ tests for adaptive). Example 6.4 below shows that this is not necessarily true for general correlation and that the gap between adaptive and non-adaptive methods can be large, highlighting that the problem may be fundamentally more challenging than its counterpart in independent group testing settings.
Claim 7.2.
Any non-adaptive algorithm with probability error at most $\frac{k}{n},k\in\{0\}\cup\mathbb{N}$, needs to do at least $n-k-1$ tests in Example 6.4.
Proof.
Suppose we have the additional information that either $e_{i-1}$ or $e_{i}$ is the target edge, $2\leq n$. Between these two cases, only node $i$ has a different state. In order to correctly decide which one is the target edge, there must be a test where $i$ is the smallest index of the node that the test contains, otherwise, no algorithm can do better than random guessing. There is $2/n$ probability that $e_{i}$ or $e_{i-1}$ is sampled, and if no test contains $i$ as its smallest index, with probability at most $1/2$ it can recover the correct edge, hence incurring an error of $1/n$. There are $n-1$ candidates for the smallest index, and by not having one of them there would be an additional $1/n$ error. hence by missing $k$ of them, we will have a probability $k/n$ of error. ∎

### v2

@@ 7 Towards Non-adaptive Group Testing
We have studied adaptive group testing so far. In Section 2 we proposed SNAGT as a model in between adaptive and non-adaptive testing. To recall, in SNAGT, tests are designed beforehand and do not depend on prior test results, which is similar to non-adaptive. But, similar to adaptive, the tests are done sequentially and we know the result of a test just after performing it. The results of the previous tests can be used to determine if the tests should stop or continue. We first propose a SNAGT strategy and analyze its performance. Then we discuss how correlation between states of different nodes complicates the design of fully non-adaptive algorithms.
We first obtain a SNAGT algorithm for building upon a non adaptive algorithm designed in [GLS22] for $d$- regular hypergraphs. The testing strategy in [GLS22] is random and each node of $V$ is picked to be in a test with probability $1/d$ independent of others. Every such test will be referred to as a random test in this section henceforth. Such tests are repeated and the selections in different tests are independent. After a predetermined number of tests, once the results are obtained, the edges that are not consistent with the result of any test, ie, those which can not be the target edge given the result, are removed. Specifically, if any test is positive then edges which do not have at least one node among the nodes tested in that test are removed; if the test is negative, edges which have at least one node tested in it are removed. It has been shown in [GLS22] that after $O(d(\log|E|+\alpha))$ tests, once the results are processed, the hypergraph contains only the target edge with probability $1-e^{-\alpha}$. Setting $e^{-\alpha}=1/(n)^{2}$ we get $\alpha=2\log n$. Thus, after $O(d(\log|E|+\log n))$ tests, the hypergraph contains only the target edge with probability $1-(1/n^{2})$ which converges to $1$ as $n\rightarrow\infty.$ Now considering that nodes are selected randomly for tests exactly as in [GLS22], but that results of a test are available right after the test, we explore if the same probability of error can be obtained through a lower number of tests by utilizing the 1) results to potentially stop earlier (SNAGT framework) and 2) knowledge of probability distribution on edges specifying the probability that any given edge is a target edge.
The core idea is as follows. We partition the edges into subgraphs so that the edge probabilities in the same subgraph are close to each other (ie, within a factor of $2$ of each other). This way, each subgraph constitutes a new group testing problem with nearly uniform probabilities where random testing, as discussed in the previous paragraph, is efficient. This idea is similar to the non-adaptive construction in [LCH+14] with two major differences: (i) In [LCH+14], nodes are partitioned, leading to several smaller subproblems. In our settings, edges are partitioned and the resulting subproblems are not smaller problems– each subproblem is still a hypergraph on $n$ vertices. (ii) the construction in [LCH+14] is non-adaptive, and finding the set of infected nodes reduces to finding a set of infected nodes in each of the smaller subproblems. In our setting, a fully non-adaptive strategy requires one to search for the infected set in all of the subgraphs with a minimum number of tests that is governed by the worst-case subproblem and not the average. Nonetheless, a semi-non-adaptive strategy is more powerful because the decoder can perform sequential posterior update after each test result to find the infected edge among all the subproblems. Here, it turns out that the minimum number of tests is governed by the average performance of the involved subproblems.
Note that a subgraph can have at most one target edge (as there is only one target edge in the system). Since the result is obtained right after a test, the result can be processed to eliminate edges that are not target edges. The target edge will never be eliminated as it will be consistent with the result of every test. We refer to the subgraph with the target edge as the target subgraph, the rest as non-target subgraphs. As edges that are not target edges are being progressively eliminated, and the target edge is never eliminated, with a high probability, after a certain number of tests, the target subgraph will have only the target edge remaining. So, any subgraph with only one remaining edge is shortlisted as a candidate subgraph for having the target edge, each such remaining edge is a candidate for being the target edge. But non-target subgraphs may also be shortlisted as candidates per the above criterion, as they may be reduced to only one edge through the same process of elimination. The number of candidates is at most the number of subgraphs. As tests continue, some of the non-target subgraphs would become empty, because of elimination of their remaining edge (which is not a target edge). If the number of tests exceed a certain threshold, with a high probability, all non-target candidate subgraphs become empty; then the non-empty candidate, if any, can be considered the target subgraph and its edge outputted as the target edge. If the number of subgraphs is not too large then the number of candidates won’t be either and the number of tests by which only the target subgraph survives (with a high probability) will not be too large either. This is indeed the case if edges with probabilities within a factor of two of each other are grouped together into the same subgraph. But again if the number of subgraphs is small, at least some of them will have a large number of edges increasing the number of tests needed to render them as candidates. But, our partitioning ensures that the edges with high probabilities of being target edges are in subgraphs with few edges each, as there can only be a limited number of high probability edges and only edges with similar probability values are grouped together. And, a subgraph with few edges will be reduced to one edge in a few tests and therefore becomes a candidate in a few tests. In contrast, the subgraphs where edges have low probabilities can reduce to one edge only in a relatively large number of tests as they can have large number of edges to start with. This quick identification of candidates among high probability edges reduce the overall number of tests, a phenomenon that suggests that high disparities between edge probabilities will yield bigger reductions in number of tests compared to that for the non adaptive algorithm in [GLS22]. We will analyze this phenomenon later, but first start with the details of the partition.
We partition the edges into subgraphs $G_{1},G_{2},\ldots,G_{k}$ where $G_{i}=(V,E_{i})$ and $\forall e\in E_{i}:\frac{1}{2^{i}}\leq p(e)\leq\frac{1}{2^{i-1}}$. Denote $m_{i}=|E_{i}|$ as the number of edges in $G_{i}$, and $\mathbb{P}(G_{i})=\sum_{e\in E_{i}}p(e)$ the probability that the target edge is in $G_{i}$. Clearly, $p(e)$ for edges in subgraph $G_{i}$ decreases with increase in $i$. For $i>n\log n$, this probability becomes vanishingly small: $p(e)<2^{-(n\log n)}$. Since there are at most $2^{n}$ edges overall, $m_{i}\leq 2^{n}$ for each $i$. Thus, for $i>n\log n$, $\mathbb{P}(G_{i})\leq\frac{2^{n}}{2^{n\log n}}$, which converges to 00 as $n\rightarrow\infty.$ We therefore consider that $k\leq n\log n.$ Since $\mathbb{P}(G_{i})\leq 1$, and $p(e)\geq\frac{1}{2^{i}}$ for each $i$, $m_{i}\leq 2^{i}$ for each $i$. Since $p(e)\leq 1/2^{i-1}$ for each $i,$ so $m_{i}\leq 2/p(e)$ for all $e\in E_{i}$. Thus, intuitively, size of a group increases with decrease in edge probabilities in the group; as such edges with highest $p(e)$ will be in subgraphs with fewest edges.
Utilizing the above partition of edges, we design a SNAGT algorithm for $d$-regular hypergraphs next. The principal challenge is in determining which candidate subgraph is the target subgraph in a reasonable number of tests. 1) We can not determine whether a candidate subgraph is the target subgraph from the timing of its candidature. This is because, as mentioned before, non-target subgraphs may also become candidates, either before or after the target subgraph becomes. Even a meaningful upper bound on the timing of candidature of the target subgraph is not known apriori. Note that it follows from [GLS22, Theorem 4], that if $G_{i}$ is the target subgraph, after $d(\log m_{i}+\alpha)$ random tests, $G_{i}$ will have only one edge left, the target edge, with a probability at least $1-e^{-\alpha}$, which is close to $1$ for large $\alpha.$ But, we do not know $i$ apriori, hence we do not know $m_{i}$ either. 2) For regular hypergraphs, testing the complement of any edge will reveal if it is the target edge or otherwise. But, the SNAGT framework does not allow us to specifically test the remaining edge or its complement, for the candidate subgraphs. This is because a subgraph becomes a candidate only as a result of outcome of tests, so testing the remaining edge in it or its complement based on the criterion that it is the candidate, would be using the test results to determine who to test. Incidentally, even if we could rule out candidate subgraphs by testing the complement of their remaining edge, we would need up to $k\leq n\log n$ tests, which is very high. 3) As mentioned before, additional random tests can determine, with a high probability, if a candidate subgraph is the target subgraph, by eliminating its remaining edge if it is not. One can in fact prove that the probability that an edge that is not the target edge can survive $d\Omega(1)$ random tests, approaches 00 as $n\rightarrow\infty.$ But, deploying up to $d\Omega(1)$ random tests after each subgraph becomes a candidate may need up to $(n\log n)(d\Omega(1))$ tests, which is $\omega(dn)$, and the maximum number of tests that can be possibly needed by the non adaptive algorithm in [GLS22] is $O(dn).$
We design an algorithm to realize the core idea while overcoming the above challenges. We consider $SG$ as a set of candidate target subgraphs. It is initialized as empty, and candidates are progressively added to $SG$ as they are identified. At any given time, $|SG|\leq k\leq n\log n.$ One can prove that up to $k$ non-target subgraphs empty out by $10d\log n$ random tests with a probability that approaches 00 as $n\rightarrow\infty$; while the target subgraph never empties. Accordingly, we consider the first candidate to enter $SG$ and survive $10d\log n$ further random tests after entering $SG$, as the target subgraph, and its sole remaining edge as the target edge. Thus, although non-target subgraphs can become candidates before the target subgraph, they would likely empty out in $10d\log n$ tests after they become candidates, and therefore not be deemed the target subgraph. The overall algorithm proceeds as follows.
  1. 1.
    Set $SG=\phi$ and $\text{time}(G_{j})=0$ for $1\leq j\leq n\log n$. If at any point, the total number of tests exceeds $2dn$, the algorithm outputs error. /*Here, the set $SG$ is the set of candidate subgraphs, ie, those with exactly one edge, and the set is updated after every test, $\text{time}(G_{j})$ is the number of random tests $G_{j}$ survives in $SG$. */
  2. 2.
    While $\nexists i:\text{time}(G_{i})\geq 10d\log n$: perform a random test, update the edges. For each $j\in SG$, update $\text{time}(G_{j})$ as $\text{time}(G_{j})+1$. Then, for all $k,l$ such that $|E_{k}|=1$, $|E_{l}|=0$, $SG=(SG\cup G_{k})\setminus G_{l}$
  3. 3.
    From the set $\{G_{i}|\text{time}(G_{i})>10d\log n\}$, choose a $G_{i}$ at random and output its single edge as the target edge.
Note that conforming to the SNAGT framework, the test designs are independent of previous test results (the nodes to be tested are always selected randomly from $V$ with the same probability distribution independent of the entire past and $V$ is not altered throughout the algorithm).
We first prove that the above algorithm returns the target edge with a probability that approaches $1$ as $n\rightarrow\infty.$ Note that once the target subgraph joins $SG$ it will never be removed from it. So the only ways the algorithm does not return a target edge is when at least one of the following two events happen: 1) the target subgraph does not enter $SG$ by $2d(n-5\log n)$ tests 2) at least one non-target subgraph that enters $SG$ before or at the same test as the target subgraph remains candidate for $10d\log n$ tests. Note that the first event implies that either the target subgraph does not become a candidate by the end of all tests or $10d\log n$ tests are not conducted between when it becomes candidate and all tests end. We refer to these events as $err_{1},err_{2}$ and show that their respective probabilities $\delta_{n}^{1},\delta_{n}^{2}$ both approach 00 as $n\rightarrow 1$. The claim made at the start of the paragraph therefore follows due to union bound.
First, we upper bound the probability of $err_{1}$ given that $G_{i}$ is the target subgraph. First note that $m_{i}\leq 2^{n}$ for each $i.$ Thus, $d\log m_{i}+2d\log n<2d(n-5\log n)$, for all large $n$. Thus, 1) the algorithm conducts at least $d\log m_{i}+2d\log n$ tests, and 2) a necessary condition for $err_{1}$ to happen is that one or more edges other than the target edge in $G_{i}$, survive $d\log m_{i}+2d\log n$ tests. Now, following the analysis in [GLS22], the probability that one or more edges other than the target edge in $G_{i}$, survive $d\log m_{i}+2d\log n$ tests, is at most $1/n^{2}$. Since this upper bound does not depend on $i$, it also upper bounds the unconditional probability of $err_{1}.$ Thus, $\delta_{n}^{1}\leq 1/n^{2}.$
We now upper bound the probability of $err_{2}.$ We will show that the probability that a given non target edge, say $e$, survives $10d\log n$ tests is at most $e^{-2.5\log n}$. Note that fewer than $k$ non-target subgraphs enter $SG$ before or at the same time as the target subgraph. Thus, together the non-target candidate subgraphs have fewer than $n\log n$ edges, none of which is the target edge. By union bound, the probability that at least one of them survives the $10d\log n$ tests is upper bounded by $n(\log n)e^{-2.5logn}\leq 1/n^{1.4}.$ Thus, $\delta_{n}^{2}\leq(1/n^{1.4})$, and the result follows. We now prove the claim at the start of the paragraph. Note that $e$ must have at least one negative node and it is detected as negative if it is selected in a test and none of the positive nodes are selected for that test. There are $d$ positive nodes overall since the target edge has $d$ nodes in a $d-$ regular hypergraph. Thus the probability that the negative node is detected in a test is $(1/d)(1-(1/d))^{d}.$ Thus in a random test $e$ is ruled out as a target edge with a probability at least $(1/d)(1-(1/d))^{d}.$ The probability that $e$ survives after $10d\log n$ tests is therefore at most
$(1-(1/d)(1-1/d)^{d})^{10d\log n}\leq e^{-10(1-1/d)^{d}\log n},$
since $1-x\leq e^{-x}$ with $(1/d)(1-(1/d))^{d}$ as $x.$ Now, noting that $(1-1/d)^{d}\geq 1/4$ for $d\geq 2$, $e^{-10(1-1/d)^{d}\log n}\leq e^{-2.5\log n}$. The claim has been proven then.
We now seek to compute the expected number of tests for the above algorithm. We use the following insight. We have shown in the previous paragraph that the target edge is returned with probability $1-\delta_{n}^{1}-\delta_{n}^{2}$. We now upper bound the number of tests in which it is returned with this probability. As noted earlier, given that $G_{i}$ is the target subgraph, after $d\log m_{i}+2d\log n$ tests, $G_{i}$ enters $SG$ (ie becomes candidate) with probability $1-\delta_{n}^{1}$. Once it enters $SG$, the algorithm conducts exactly $10d\log n$ subsequent tests, because $G_{i}$ survives all of them and the algorithm terminates after at least one candidate survives $10d\log n$ tests. So with probability more than $1-1/n^{2}$, it performs $d\log m_{i}+2d\log n+10d\log n$ tests conditioned on $G_{i}$ being the target subgraph, which happens with probability $\mathbb{P}(G_{i}).$ Conditioned on $G_{i}$ being the target subgraph, with probability at most $\delta_{n}^{1}\leq 1/n^{2}$, $G_{i}$ does not enter $SG$ by $d\log m_{i}+2d\log n$ tests, but even so the algorithm can use at most $2dn$ tests. Thus, the expected number of tests, $\mathbb{E}[T],$ is at most
$\sum_{i}\mathbb{P}(G_{i})[(1-\delta_{n}^{1})(d\log(m_{i})+2d\log n+10d\log n)+(1/n^{2}*2dn)]$
$\leq d\sum_{i}\left[(\sum_{e\in G_{i}}p(e))\log(m_{i})\right]+12d\log n+O(d/n).$
Recall that $m_{i}\leq 2/p(e)$ for all $e\in E_{i}$. Hence the RHS is
$\leq d\sum_{i}\sum_{e\in G_{i}}p(e)\log(2/p(e))+12d\log n+O(d/n)$
$=d\sum_{e\in E}p(e)(1+\log(1/p(e)))+12d\log n+O(d/n)$
$=d+dH(X)+12d\log n+O(d/n).$ (10)
So, we get $O(dH(X)+d\log n)$ tests in expectation.
We have therefore proven the following theorem for the SNAGT framework for a $d-$regular hypergraph (considering that $u=d$ for any such hypergraph):
@@
Consider a graph $G$ with $I$ infected nodes in it. Let $\Pr(I\geq u)$ be $o(1/n)$ for $u\geq 2$. There is an algorithm that performs $O(uH(X)+u\log n)$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
In Appendix C we prove the above theorem when the hypergraph is not regular.
Now, [GLS22] provides a non-adaptive combinatorial group-testing algorithm that finds the target edge in a $d$-regular hypergraph in $O(d(\log|E|+\log n))$ tests, with probability at least $1-(1/n^{2}).$ Our algorithm is semi non-adaptive, and there is a probability distribution on edges for being the target edge. We obtain the same result as [GLS22] in $O(d(H(X)+\log n))$ tests in the SNAGT framework. When the probability over the edges is uniform, which maximizes the entropy, the number of tests for our algorithm is order-wise the same as [GLS22]. It is already known that there is a close connection between combinatorial group testing and probabilistic group testing when the distribution is uniform. When the distribution is not uniform, $H(X)$ can be significantly smaller than $\log|E|$. Particularly, the testing is significantly improved when $|E|$ is $\Omega(n)$ and the distribution over the edges is far from uniform, ie, $p(e)$ is large for some edges. Note that $m_{i}$ is small for a subgraph $i$ in which $p(e)$ is large. If such a subgraph is the target subgraph, it becomes a candidate with high probability in $O(d(\log m_{i}+\log n))$ tests and the algorithm terminates after conducting $10d\log n$ additional tests. The total number of tests is small then if an edge with high probability of being a target edge is the target edge, because $m_{i}$ is small for subgraphs containing such edges.
The SNAGT algorithm needs larger, by a multiplicative factor of $d$ or $\log n$, number of tests than the adaptive algorithm. But it is easier to execute. Not only does it not need results of any test to determine who to test next, it does not need knowledge of exact probability that an edge is a target edge; it only needs to know for which $i$ is this probability between $2^{-i}$ and $2^{-(i-1)}.$
Finally, in Theorem 7.1 we have assumed that the number of infected nodes is less than an upper bound $u$, equivalently $|e^{*}|\leq u$, with high probability. This would for example hold if the number of infected nodes is concentrated around its mean, $\mu$, then $u=\mu+o(\mu)$. Incidentally, when the testing is done adaptively, we get the best results when the number of infections is concentrated around its mean $\mu$ (eg, Corollary 4.12, Section 4). Also, the testing algorithm, which is described in Appendix C, and which has been used to prove Theorem 7.1, does not need knowledge of $\mu$, but only of a stochastic upper bound on the number of infected nodes. Thus, the algorithm is intrinsically robust to errors in the estimation of the system parameters. Note that the upper bound on the expected number of tests for this Algorithm decreases with decrease in the stochastic upper bound on the number of infected nodes. Thus, if a tighter upper bound on the number of infected nodes can be estimated, then the guarantee on the performance of the Algorithm improves. Clearly, $u\geq\mu.$ Thus, the guarantee on the expected number of tests decreases at most to $O(\mu H(X)+\mu\log n)$.
We have not been able to generalize the above technique to get general upper bounds for non-adaptive group testing. In our algorithm, the encoding stage (i.e., test design) is straightforward, consisting simply of random tests. However, due to the nature of our model, the decoding is more sophisticated. Specifically, it updates posterior probabilities in real time, tracks potentially candidate subgraphs, and strategically determines when to halt the algorithm. These capabilities distinguish our semi-non-adaptive approach from the non-adaptive setting, where real-time decoding is not possible.
Non-adaptive group testing has been extensively studied in the literature for independent group testing [LCH+14, WGG23, PR08], and typically, state of the art non-adaptive methods differ from the adaptive counterparts only by $O(\log n)$ or $O(d)$ (multiplicative factor). For instance, $O(d^{2}\log n)$[DHH00] is needed in the classic non-adaptive combinatorial setting (compared to $O(d\log n)$ tests for adaptive) or $O(\log n(H(X)+\mu))$[LCH+14] in non-adaptive independent group testing (compared to $O(H(X)+\mu)$ tests for adaptive). Example 6.6 below shows that this is not necessarily true when states of nodes are correlated and that the gap between adaptive and non-adaptive methods can be large, highlighting that the problem may be fundamentally more challenging than its counterpart in independent group testing settings.
Claim 7.2.
Any non-adaptive algorithm with probability error at most $\frac{k}{2n},k\in\{0\}\cup\mathbb{N}$, needs to do at least $n-k-1$ tests in Example 6.6.
@@
Consider a non adaptive testing policy which conducts $l$ tests, the nodes tested in each test are given in $P=\{T_{1},T_{2},\ldots,T_{l}\}$ where $T_{i}\subseteq V$ is the set of nodes tested in the $i$th test. We argue that without loss of generality we can consider that each $T_{i}$ consists of only one node, otherwise consider that some $T_{i}$ consists of nodes $p,q$, where $q>p.$ Then if $q$ is positive, so is $p$, and removing $q$ from $T_{i}$ will yield the same result. Next, if $q$ is negative, removing $q$ from $T_{i}$ will yield the same result. Thus, $T_{i}$ will yield the same result if only its left-most node is tested. Thus, we consider that $P=\{i_{1},i_{2},\ldots,i_{l}\}.$ Note that $l\leq n-1$, as node $1$ is positive, since otherwise all nodes are negative and there is no target edge, and thus, it is redundant to test node $1$
If more than one edge is consistent with the test results, the algorithm can do no better than to randomly (and uniformly) choose one of them as the target edge. Let $i\not\in P$, $i\geq 2.$ Note that $e_{i-1}\subset e_{i}$, and $e_{i}\setminus e_{i-1}=\{i\}.$ A necessary condition for $e_{i-1}$ to be determined as the target edge is that $e_{i}$ be ruled out while $e_{i-1}$ is not ruled out. This can happen only if $i$ tests negative. If $e_{i-1}$ is the target edge, it can not be ruled out by any test, but then since $i$ is not tested, the status of node $i$ can not be determined, hence $e_{i}$ can not be ruled out either.
Since $2\leq i\leq n$, we considered edges in $\{e_{1},e_{2},\ldots,e_{n-1}\}$ in the above argument. The above argument shows that at most for $l$ edges, among $e_{1},\ldots,e_{n-1}$, the algorithm can have zero error given that any of these is the target edge, and for the rest $n-1-l$ edges there is at least a probability of $1/2$ for the algorithm to output a wrong edge given that one of them is the target edge. Given that every edge is equally likely to be the target edge in Example 6.6, the probability of error in this case is $\frac{n-1-l}{2n}$. Substituting $k$ with $n-l$ will prove the result.